The true average weight of a bag should be 50 kg. 36 bags are weighed obtaining an average of 49.5 kg with a deviation of 1.2 kg. Do a hypothesis test, with 95% confidence, to verify if the content of the bags is different from 50 kg.

A) z = -2.5 <-1.96; Ho is rejected
B) z = -2.5 <-1,645; Ho is rejected
C) z = -2.5 <-1.96; Ho is not rejected
D) z = -2.5 <-1,645; Ho is not rejected
Answer #1

Test statistic = z

= (\bar x - \mu ) / \sigma    / \sqrt n

= (49.5 - 50) / 1.2 / \sqrt 36

=  -2.5

Test statistic =  -2.5

A) z = -2.5 <-1.96; Ho is rejected